OpenStudy (anonymous):

Choose the best description for the real number 4 point 6 repeating. Irrational, because it is a terminating decimal Irrational, because it is a repeating decimal Rational, because it is a terminating decimal Rational, because it is a repeating decimal

OpenStudy (mathmate):

Hint: All repeating decimal numbers can be converted to a fraction. If you look up the definition for rational numbers, you will find the correct answer.

OpenStudy (lalipapa123):

.................

OpenStudy (mathmate):

For example, 0.66666... equals (exactly) 2/3 1.01212121212.... = 1+12/990=1002/990=167/165

OpenStudy (anonymous):

So...D? Sorry, i haven't done this in a long time so i forgot most of this..

OpenStudy (mathmate):

Look up the definition of rational numbers, and together with the fact that all repeating decimals can be converted to a fraction, you will be confident of your answer. Here's a nice link for the definition of rational numbers, and it is short but explicative: https://www.mathsisfun.com/definitions/rational-number.html